All Systems Operational
Norce Merchant Center [Storm] ? Operational
90 days ago
99.99 % uptime
Today
Storm API ? Operational
90 days ago
99.96 % uptime
Today
Storm Connect ? Operational
Storm Query ? Operational
Storm Event ? Operational
Storm Internals ? Operational
Storm Imageserver ? Operational
Storm Stage ? Operational
Storm Export Feeds ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Storm API Response times
Fetching
Storm Admin
Fetching
Storm API Uptime
Fetching
Past Incidents
Nov 30, 2022

No incidents reported today.

Nov 29, 2022

No incidents reported.

Nov 28, 2022

No incidents reported.

Nov 27, 2022

No incidents reported.

Nov 26, 2022
Resolved - Order throughtput was increased and the order queue was caught up at ca 01:00.
Nov 26, 09:39 CET
Monitoring - Due to very high Order throughput, we're experiencing delays in order delivery. Expect delays for payed basket to be transformed to orders and delivered to customer ERP systems. We're scaling up to be able to deliver more orders per minute.
Nov 25, 22:22 CET
Nov 25, 2022
Resolved - Storm's Data Center Provider has confirmed the network error. It was caused by a change in a central switch in the backup data center which had an unexpected effect of disrupting network connections in the primary data center. The change is rolled back and will not be attempted again.
Nov 25, 15:23 CET
Investigating - We've experience another internet connectivity issue, and are working closely with our data center provider to prevent the problem from occurring again.
Nov 25, 15:03 CET
Resolved - The problem has been resolved. Our service provider is investigating the internet connectivity problem further, the problem affected the entire data center and the internet connectivity was lost for 1 minute 35 seconds at 14:15.
Nov 25, 14:41 CET
Update - We are continuing to investigate this issue.
Nov 25, 14:38 CET
Investigating - We've experienced internet connectivity loss to Storm datacenter, we were down 14:15-14:16. We're in contact with our Internet Service Provider to determine the extent of the problem. We're currently up.
Nov 25, 14:23 CET
Nov 24, 2022

No incidents reported.

Nov 23, 2022

No incidents reported.

Nov 22, 2022

No incidents reported.

Nov 21, 2022
Resolved - We're back to normal price update rates and delays.
Nov 21, 23:02 CET
Monitoring - We're currently having substantially more price updates than normal high load, which means that price updates takes longer time to be activated. We don't experience any problems, but expect longer delays than normal.
Nov 21, 20:40 CET
Nov 20, 2022
Resolved - The queue is almost processed now with new order service instances. Issue resolved.
Nov 20, 13:52 CET
Monitoring - Order service have received more orders than it can handle. We have scaled out service to five more instances to handle order load and monitoring performance. You may experience lag for a while until order queue has been processed.
Nov 20, 13:45 CET
Nov 19, 2022

No incidents reported.

Nov 18, 2022
Resolved - The problem has been resolved. Parallelism has been increased and the queue is decreasing rapidly and will be caught up shortly.
Nov 18, 14:37 CET
Update - After analysis we've discovered that orders are being created and sent to Customer ERP systems, but are currently delayed ca one hour. We're working on increasing throughput.
Nov 18, 12:42 CET
Investigating - We're currently having problems with payed baskets not being completed to orders, and therefore orders are not transmitted to customer ERP. We are investigating the issue.
Nov 18, 12:20 CET
Nov 17, 2022

No incidents reported.

Nov 16, 2022
Resolved - We continue working on the issue, the performance is nearly back to normal. The problem is related to recurring database maintenance which has not been executed correctly and has caused the performance issue. We need to perform further analysis and will continue to monitor to ensure the correct actions have been taken. We plan to publish a post-mortem during Thursday 17th.
Nov 16, 18:45 CET
Update - We have to a certain extent restored services, we continue to work on the issue.
Nov 16, 16:56 CET
Update - StormApi is currently unavailable, we continue to work on resolving the problem.
Nov 16, 16:35 CET
Update - We continue to investigate the issue, we have currently no prognosis to when the problem may be resolved.
Nov 16, 16:07 CET
Investigating - We're currently having, and have had intermittent problems, with performance in StormApi, we're investigating the problem.
Nov 16, 15:26 CET